OpenStudy (anonymous):

ill give medal:) Which function below has the largest slope? f(x) x f(x) 2 4 4 12 6 20 g(x) g(x)=10x - 4 h(x) Lance gains 1 level each time he plays his favorite video game. He started at level 10. f(x) g(x) h(x) The functions all have the same slope

OpenStudy (anonymous):

slope=m ; m for f(x)=4 n for g(x)=10..

OpenStudy (anonymous):

but can u plz explain more about h(x)..

OpenStudy (anonymous):

that's all it gives me

OpenStudy (anonymous):

both m are calculated as three ordered pairs mentioned in question.. m= y2-y1/ x2-x1

OpenStudy (anonymous):

so in question 3 pairs for f(X) r mentioned, take their difference n get m = 4

OpenStudy (anonymous):

then find g(x) by putting values of x in equation of g(x) then u'll get 3 pairs again.. again calculate m u get m=10..

OpenStudy (anonymous):

im confused.. so the answer is that g(x) has the greatest slope?

OpenStudy (anonymous):

but for h(X) no details are given .. but if h(x) is started at 10 and every time he gain 1 level means its greater than 10.. rite???

OpenStudy (anonymous):

no need 2 cnfuse here dear.. its simple...

OpenStudy (anonymous):

ok truu. ur right

OpenStudy (anonymous):

Which function below has the smallest slope? f(x) f(x) = 2x - 3 g(x) graph going through (0, negative 5) and (4, 3)

OpenStudy (anonymous):

f(x) has smallest slope...

OpenStudy (anonymous):

frst u hav 2 find the m value for 2 values of g(x) using the same formula u'll get m= 3-(-5) divided by 4-0 so m=8/4=2

OpenStudy (anonymous):

then use the same values of x n put in the above equation for f(x) u'll get (0,3) &(4,5) find m=1/2 hence f(x) has smallest slope

OpenStudy (anonymous):

does f(x) have a smaller slope than h(x) x h(x) -1 3 1 7 3 11

OpenStudy (anonymous):

z this continuation???

OpenStudy (anonymous):

m for the mentioned h(x)= 4/2=2

OpenStudy (anonymous):

if it'f the continuation then h(x) n g(x) have m=2 but f(x)'m=1/2

OpenStudy (anonymous):

rite?